# Grounding Foundation Models through Federated Transfer Learning: A General Framework

### 10 Dec 2023 v5

@@ 5.4. Privacy and Backdoor Attacks
Privacy attacks focus on inferring private data from knowledge at the representation level. More specifically, TextObfuscator (Zhou et al., 2023), RAPT (Li et al., 2023b) and Xu et al. (2023c) investigated model inversion attacks in federated split learning scenarios, in which the server aims to reconstruct the client’s private data based on the representations output from the client’s local model. These studies proposed representation perturbation, $d_{\chi}$-privacy, and weight permutation equivalence, respectively, to thwart model inversion attacks. The rationale behind these defenses is to reduce the correlation between private data and the representations shared with the server. As a result, the chance for the adversary to recover private data can be reduced.
@@
Regarding hetero-model transfer, Li et al. (2023d) proposed BD-FMFL, a backdoor attack following the framework proposed in GPT-FL (Zhang et al., 2023a). BD-FMFL works as follows: the attacker inserts backdoors into a Language Model with Large Memory (LLM) by leveraging in-context learning. Subsequently, the server utilizes the backdoored LLM to generate prompts, which are then used to synthesize data using other FMs. The server proceeds to train a global model based on the synthetic data, thereby transferring the backdoor to the global model. Once accomplished, the backdoored global model is fine-tuned using private datasets provided by clients within the standard Federated Learning (FL) framework. Another avenue of research, explored by BadPrompt (Cai et al., 2022) and PPT (Du et al., 2022b), focuses on prompt-based backdoor attacks, in which the attacker injects backdoors into the prompt model of an FM and subsequently releases the poisoned prompt model to the public. Upon a client downloading and utilizing the compromised prompt model for downstream tasks, the server gains the ability to activate the backdoor. BadPrompt suggests that methods such as Fine-Pruning (FP) and knowledge distillation could potentially defend against prompt-based backdoor attacks.
6. Domain Knowledge for Augmenting Foundation Models
In this section, we overview FTL-FM methods that fall into setting 2⃝ (see Definition 3.1), the objective of which is to augment the FM with industry-level and domain-specific knowledge (e.g., healthcare, medicine, finance, banking, and law) transferred from domain-specific clients.
@@ 6. Domain Knowledge for Augmenting Foundation Models
FedOST (WeBank, [n. d.]) $\bigcirc \bigcirc{}^{\dagger}$ Adapter Emulator of $\mathcal{W}^{S}_{\text{FM}}$, Initial adapter SA CP
@@
The key to augmenting FMs to be more knowledgeable, accurate, and reliable lies in fine-tuning these FMs using high-quality domain-specific datasets. These high-quality datasets may contain private and sensitive information and are often scattered among isolated entities (devices and companies). FTL-FM methods enable the server of FM and clients with domain knowledge to augment the FM collaboratively while preserving the privacy of clients’ knowledge. In literature, representation-level knowledge transfer and model-level knowledge transfer are the main research directions explored to augment an FM hosted by a server with knowledge transferred from domain-specific clients.
@@ 6.1. Domain knowledge transferred to FM through representation-level knowledge transfer
Toward this end, CreamFL (Yu et al., 2023b) was proposed to learn an FM (i.e., $\text{BERT}_{\text{base}}$) by transferring representations from diverse clients that own uni- and multi-modal data. In CreamFL, each DM and the FM are split into top and bottom parts. Upon receiving global representations, each client first trains its DM based on its local data and global representations. Then, each client generates local representations using its domain bottom model based on a public dataset and sends local representations to the server. The server, in turn, aggregates local representations sent from clients with global representations and trains the foundation bottom model through knowledge distillation based on aggregated representations. Next, the server sends global representations generated from the FM to clients for the next round of training. The training process of CreamFL is illustrated in Figure 8(a). CreamFL generally adheres to the objective formulated in Eq.(3) with $K>1$, $\sum^{K}_{k=1}p^{C_{k}}=1$, $p^{S}>0$, and $\eta=0$. The specific form of the loss $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ for transferring representations generated by $\mathcal{W}^{C_{k}}$ to train $\mathcal{W}^{S}_{\text{FM}}$ is formulated in Eq.(11).
@@ 6.2. Domain knowledge transferred to FM through model-level knowledge transfer
To alleviate these limitations, Fan et al. (Fan et al., 2023) proposed FedOST, a federated learning version of Offsite-Tuning. In FedOST, the server first distributes the emulator and two adapters of its FM to all clients at the beginning of FL. Then, each client fine-tunes the two adapters on its private data with the assistance of the emulator. Next, all clients send their adapters to the server, which aggregates received adapters through secure aggregation (Bonawitz et al., 2016) and plugs the aggregated adapter into the FM. FedOST protects the privacy of clients’ private data and can build a fine-tuned FM with enhanced generalization capability. With similar motivations as FedOST, Kuang et al. integrated Offsite-Tuning into FederatedScope-LLM (Kuang et al., 2023b) and termed it FedOT. FedPEAT (Chua et al., 2023) generalized the Offsite-Tuning approach to federated learning with multiple communication rounds and proposed an adaptive control mechanism to facilitate the adoption of FedPEAT in a dynamic environment. CEFHRI (Khalid et al., 2023) adopted Offsite-Tuning to the human-robot interaction domain.
Similar to Offsite-Tuning, FedOST, FedOT, and FedPEAT adhere to the objective formulated in Eq.(3) but with $K>1$, $\sum^{K}_{k=1}p^{C_{k}}=1$, and $p^{S}=0$. FedOST has $\eta>0$ since it explicitly leverages secure aggregation to protect data privacy. In general, the training process of the three methods can be illustrated in Figure 8(b).
@@ 6.3. Privacy and Backdoor Attacks
7. Co-optimize Foundation Models and Domain Models
The FTL-FM settings we discussed in Section 5 and Section 6 focus on adapting general knowledge of the FM to optimize DMs, and transferring domain-specific knowledge of DMs to optimize FM, respectively. In this section, we explore the setting 3⃝, which endeavors to mutually optimize FMs and DMs hosted by different parties (i.e., the server and clients).
@@ 7. Co-optimize Foundation Models and Domain Models
FedLLM on Edge (Woisetschläger et al., 2023) $\mathcal{W}^{C}$ Global model
@@
FedDAT (Chen et al., 2023b) Dual-adapter Global dual-adapter
FedIT (Zhang et al., 2023b) LoRA model Global LoRA model
@@ 7.1. Co-optimize FMs and DMs through model-level knowledge transfer
    The server owns an FM, while clients own DMs. Clients’ DMs can optionally be initialized with FMs. The objective of federated learning is to augment the server’s FM while training (or fine-tuning) clients’ DMs. As a result, the server ends up with an FM augmented with domain-specific knowledge, and each client with a DM enhanced with general knowledge of the FM.
@@
    The server serves as an aggregation function while clients’ domain models are initialized with FMs. The objective of federated learning is to fine-tune clients’ DMs through federated learning. Each client ends up with a DM enhanced with the domain knowledge of all clients.
Figure 9. Co-tuning server’s FM and clients’ FM through transferring model-level knowledge among participating parties (both server and clients).
For the first scenario, FATE-LLM (Fan et al., 2023) proposed a FedCoLLM (Federated Co-tuning LLM) framework (Fan et al., 2023) to co-tune the FM of the server and DMs of clients. Figure 9(a) overviews the workflow of FedCoLLM. In FedCoLLM, clients are initialized with LLaMA -7B (Touvron et al., 2023), and they send their LLaMa-7B models to the server with secure aggregation (Bonawitz et al., 2016). On the server side, the server distills the knowledge between its LLaMa-65B model and the aggregated LLaMA-7B model to co-evolve models on both sides. When the clients have insufficient computing resources, the co-tuning of the FM and DMs can be achieved with the help of PEFT techniques, as illustrated in Figure 9(b). FedCoLLM correspond to the objective formulated in Eq.(4) with $K>1$, $\sum^{K}_{k=1}p^{C_{k}}=1$, $p^{S}>0$ and $\eta>0$, and the specific form of the loss $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ for co-tuning the server’s $\mathcal{W}^{S}_{\text{FM}}$ and a client $k$’s $\mathcal{W}^{C_{k}}$ is formulated in Eq.(12).
@@
For the second scenario, FedNLP (Lin et al., 2022) provided a benchmarking framework for evaluating federated learning algorithms (FedAvg, FedProx, and FedOPT) on common NLP tasks with foundation models (e.g., BART). FedNLP is computationally expensive, primarily because of the large sizes of foundation models and the complexity of NLP tasks (Cai et al., 2023b). To alleviate this issue, various methods have been proposed to leverage Parameter-Efficient Fine-Tuning (PEFT) techniques to fine-tune clients’ DMs initialized with pre-trained FMs, as illustrated in Figure 10. These methods correspond to a special case of the objective formulated in Eq.(5) and the specific form of the loss $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ for fine-tuning client $k$’s $\mathcal{W}^{C_{k}}$ with PEFT techniques is formulated in Eq.(13).
@@
Figure 10. Federated parameter-efficient-finetuning of clients’ domain FMs. Clients’ DMs initialized with FMs learn from each other through federated fine-tuning of their PEFT modules (e.g., adapter, prompt model, and LoRA model). The server acts as an aggregation function.
@@
As for now, few works following Eq.(5) explicitly consider protecting data privacy (i.e., $\eta>0$). FedPETuning (Zhang et al., 2023c), PEU+LoRA (Xu et al., 2023a), and FedPrompt (Zhao et al., 2023b) are representative ones. FedPETuning leverages four PEFT techniques, namely adapter tuning, prefix tuning, LoRA (Hu et al., 2022), and BitFit (Ben Zaken et al., 2022), to fine-tune clients’ DMs (i.e., $\text{Roberta}_{\text{base}}$) and demonstrated that federated learning combined with LoRA achieved the best privacy-preserving results. PEU+LoRA applies DP with Partial Embedding Updates (PEU) and LoRA to fine-tune clients’ DMs, achieving better privacy-utility-resource trade-offs than baselines. Zhao et al. (2023b) proposed FedPrompt that leverages federated learning combined with soft prompt tuning to fine-tune clients’ DMs (e.g., T5 (Raffel et al., 2020)), aiming to reduce communication cost while maintaining model accuracy. FedPrompt adopts local DP to protect the privacy of clients’ data, demonstrating there is a trade-off between privacy and model accuracy.
@@
Most methods following Eq.(5) focus on improving computation and communication efficiency (i.e., $\eta=0$). Chen et al. (2023b) proposed FedDAT, in which each client fine-tunes its local multi-modal model (e.g., VAuLT (Chochlakis et al., 2022)) by leveraging a Dual-Adapter Teacher (DAT) module to achieve better convergence rate than using conventional PEFT methods. Li et al. (2022b) proposed FedSplitBERT, which splits the BERT encoder into a local part and a global part and shares only a quantized version of the global part with the server for aggregation, aiming to reduce the communication cost with minimal performance loss. FedIT (Zhang et al., 2023b) was proposed to leverage federated learning combined with LoRA to efficiently utilize clients’ diverse instructions stored on local devices to improve the performance of global FM (e.g., Vicuna (Chiang et al., 2023)). Che et al. (2023) proposed a parameter-efficient prompt tuning approach with adaptive optimization, named FedPepTAO, to fine-tune clients’ large language models (e.g., LLaMA-7B) in a non-IID federated learning setting. Halbe et al. (2023) proposed HePCo to address a continual federated learning problem. In HePCo, each client is deployed with a ViT model, and all clients share only a prompt model and a classifier (on top of the ViT) with the server to save communication costs. To tackle catastrophic forgetting brought by continual learning and data heterogeneity across clients, the server trains a generator based on current and previous tasks and then distills the knowledge from the generator to the global prompt model and classifier.  Wu et al. (2023b) proposed Fed-LPFM, in which each client uses its local data and a set of private FMs to train a proxy model, and clients’ proxy models are shared with and aggregated by the server. At inference, only the proxy models are used. As a result, the inference latency can be considerably reduced and the data distribution gap between clients can be mitigated.
@@
In addition to the aforementioned works, a potentially more demanding and intriguing avenue of research involves empowering edge devices (e.g., smartphones, wearables, and Internet of Things devices) with the knowledge and capabilities of FMs, thereby greatly augmenting the intelligence of systems facilitated by these devices. Take, for instance, the case of smart home systems (King et al., 2023) enhanced by FMs, which can elevate the quality of life for individuals, or the case of surveillance systems (Yao et al., 2019) fortified by FMs, leading to heightened efficiency and accuracy in detecting moving objects such as people and traffic. To achieve these visions, the primary challenge lies in effectively and efficiently adapting FMs to edge devices that have limited memory and computing resources. Federated learning presents a promising solution to this challenge, and several studies have investigated this direction. FedAdapter (Cai et al., 2023b) was proposed to improve the computational and communication efficiency of fine-tuning local FMs (e.g., BERT) resided in devices through PEFT (i.e., adapter). To achieve this, it dynamically adjusts the depth and width of adapters during training. FeS (Cai et al., 2023a) was proposed to address the problem of data scarcity in mobile NLP applications by utilizing pseudo-labeling. It employs a curriculum pacing mechanism to gradually speed up the pseudo-labeling speed, a representational filtering mechanism to pseudo-label samples that are most beneficial to learn from, and a co-planing mechanism to control the layer depth and capacity to be trained. Furthermore, FeS fine-tunes only bias parameters of FMs (e.g., RoBERTa-large) while keeping the other parameters frozen to improve training efficiency further. FwdLLM (Xu et al., 2023b) aims to enhance the efficiency of PEFT techniques. It observed that PEFT techniques mainly alleviate network bottlenecks but do not significantly improve training efficiency as they require backpropagation across the entire model to obtain gradients. To address this challenge, FwdLLM employs a backward propagation (BP)-free method to estimate gradients during the forward pass, thereby saving substantial memory and training time. To tackle system heterogeneity, Cho et al. (2023) proposed a heterogeneous LoRA (HLoRA) approach that can apply different rank LoRA modules to devices with heterogeneous and constrained computing resources by leveraging zero-padding and truncation for the aggregation. Woisetschläger et al. (2023) conducted an in-depth study on hardware performance optimization for federated fine-tuning large language models at the edge. More specifically, It investigated the state-of-the-art deep learning accelerators as well as optimization levers for FL workloads on embedded hardware.

### 29 Dec 2023 v7

@@ 5.4. Privacy and Backdoor Attacks
Privacy attacks focus on inferring private data from knowledge at the representation level. More specifically, TextObfuscator (Zhou et al., 2023), RAPT (Li et al., 2023c) and Xu et al. (2023c) investigated model inversion attacks in federated split learning scenarios, in which the server aims to reconstruct the client’s private data based on the representations output from the client’s local model. These studies proposed representation perturbation, $d_{\chi}$-privacy, and weight permutation equivalence, respectively, to thwart model inversion attacks. The rationale behind these defenses is to reduce the correlation between private data and the representations shared with the server. As a result, the chance for the adversary to recover private data can be reduced.
@@
Regarding hetero-model transfer, Li et al. (2023e) proposed BD-FMFL, a backdoor attack following the framework proposed in GPT-FL (Zhang et al., 2023b). BD-FMFL works as follows: the attacker inserts backdoors into a Language Model with Large Memory (LLM) by leveraging in-context learning. Subsequently, the server utilizes the backdoored LLM to generate prompts, which are then used to synthesize data using other FMs. The server proceeds to train a global model based on the synthetic data, thereby transferring the backdoor to the global model. Once accomplished, the backdoored global model is fine-tuned using private datasets provided by clients within the standard Federated Learning (FL) framework. Another avenue of research, explored by BadPrompt (Cai et al., 2022) and PPT (Du et al., 2022b), focuses on prompt-based backdoor attacks, in which the attacker injects backdoors into the prompt model of an FM and subsequently releases the poisoned prompt model to the public. Upon a client downloading and utilizing the compromised prompt model for downstream tasks, the server gains the ability to activate the backdoor. BadPrompt suggests that methods such as Fine-Pruning (FP) and knowledge distillation could potentially defend against prompt-based backdoor attacks.
6. FTL-FM: Domain Knowledge for Augmenting Foundation Models
In this section, we overview FTL-FM methods that fall into setting 2⃝ (see Definition 3.1), the objective of which is to augment the FM with industry-level and domain-specific knowledge (e.g., healthcare, medicine, finance, banking, and law) transferred from clients.
@@ 6. FTL-FM: Domain Knowledge for Augmenting Foundation Models
FedOST (WeBank, [n. d.]; Fan et al., 2023) $\bigcirc \bigcirc{}^{\dagger}$ Adapter Emulator of $\mathcal{W}^{S}_{\text{FM}}$, Initial adapter SA CP
@@
The key to augmenting FMs to be more knowledgeable, accurate, and reliable lies in fine-tuning these FMs using high-quality domain-specific datasets. These high-quality datasets may contain private and sensitive information and are often scattered among isolated entities (devices and companies). FTL-FM methods enable the server of FM and clients with domain knowledge to augment the FM collaboratively while preserving the privacy of clients’ knowledge. In literature, representation-level knowledge transfer and model-level knowledge transfer are the main research directions explored to augment an FM hosted by a server with knowledge transferred from domain-specific clients. We summarize FTL-FM works that fall into this setting in Table 6, and in the following subsections, we will elaborate on these works.
@@ 6.1. Domain knowledge transferred to FM through representation-level knowledge transfer
Toward this end, CreamFL (Yu et al., 2023b) was proposed to learn an FM (i.e., $\text{BERT}_{\text{base}}$) by transferring representations from diverse clients that own uni- and multi-modal data. In CreamFL, each DM and the FM are split into top and bottom parts. Upon receiving global representations, each client first trains its DM based on its local data and global representations. Then, each client generates local representations using its domain bottom model based on a public dataset and sends local representations to the server. The server, in turn, aggregates local representations sent from clients with global representations and trains the foundation bottom model through knowledge distillation based on aggregated representations. Next, the server sends global representations generated from the FM to clients for the next round of training. The workflow of CreamFL is illustrated in Figure 10(a). CreamFL generally adheres to the objective formulated in Eq.(3) with $K>1$, $\sum^{K}_{k=1}p^{C_{k}}=1$, $p^{S}>0$, and $\eta=0$. The specific form of the loss $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ for transferring representations generated by $\mathcal{W}^{C_{k}}$ to train $\mathcal{W}^{S}_{\text{FM}}$ is formulated in Eq.(11).
@@ 6.2. Domain knowledge transferred to FM through model-level knowledge transfer
To alleviate these limitations, Fan et al. (Fan et al., 2023) proposed a federated learning version of Offsite-Tuning named FedOST. In FedOST, the server first distributes the emulator and two adapters of its FM to all clients at the beginning of FL. Then, each client fine-tunes the two adapters on its private data with the assistance of the emulator. Next, all clients send their adapters to the server, which aggregates received adapters through secure aggregation (Bonawitz et al., 2016) and plugs the aggregated adapter into the FM. FedOST protects the privacy of clients’ private data and can build a fine-tuned FM with enhanced generalization capability. With similar motivations as FedOST, Kuang et al. integrated Offsite-Tuning into FederatedScope-LLM (Kuang et al., 2023b) and termed it FedOT. FedPEAT (Chua et al., 2023) generalized the Offsite-Tuning approach to federated learning with multiple communication rounds and proposed an adaptive control mechanism to facilitate the adoption of FedPEAT in a dynamic environment. CEFHRI (Khalid et al., 2023) adopted Offsite-Tuning to the human-robot interaction domain.
Similar to Offsite-Tuning, FedOST, FedOT, and FedPEAT adhere to the objective formulated in Eq.(3) but with $K>1$, $\sum^{K}_{k=1}p^{C_{k}}=1$, and $p^{S}=0$. FedOST has $\eta>0$ since it explicitly leverages secure aggregation to protect data privacy. In general, the workflow of the three methods can be illustrated in Figure 10(b).
@@ 6.3. Privacy and Backdoor Attacks
CrossLM (Deng et al., 2023) $\mathcal{W}^{C}$ $\mathcal{W}^{C}$ enhanced by $\mathcal{W}^{S}_{\text{FM}}$
FedLLM on Edge (Woisetschläger et al., 2023a) $\mathcal{W}^{C}$ Global model
@@
FedDAT (Chen et al., 2024) Dual-adapter Global dual-adapter
FedIT (Zhang et al., 2023c) LoRA model Global LoRA model
@@
CLIP2FL (Shi et al., 2024) CLIP-enhanced $\mathcal{W}^{C}$ and gradients of local classifier Global model
@@
7. FTL-FM: Co-optimize Foundation Models and Domain Models
The FTL-FM settings we discussed in Section 5 and Section 6 focus on adapting general knowledge of the FM to optimize DMs, and transferring domain-specific knowledge of DMs to optimize FM, respectively. In this section, we explore the setting 3⃝, which endeavors to mutually optimize FMs and DMs hosted by different parties (i.e., the server and clients). We summarize FTL-FM works that fall into this setting in Table 7, and in the following subsections, we will elaborate on these works.
@@ 7.1. Co-optimize FMs and DMs through model-level knowledge transfer
    The server owns an FM and clients own DMs. Clients’ DMs can optionally be initialized with FMs. The objective of this scenario is to optimize the server’s FM as well as clients’ DMs. Consequently, the server ends up with an FM augmented with domain-specific knowledge, and each client with a DM enhanced with general knowledge of the FM.
@@
    The server serves as an aggregation function while clients’ domain models are initialized with or assisted by FMs. The objective of this scenario is to fine-tune clients’ DMs through federated learning. Each client ends up with a DM enhanced with the domain knowledge of all clients.
Figure 11. Co-optimizing server’s FM and clients’ FM through transferring model-level knowledge among participating parties (i.e., server and client). (a) illustrates the federated co-optimization of clients’ full DMs and the server’s full FM. (b) illustrates federated parameter-efficient co-optimization of clients’ DMs and the server’s FM when the clients initialize their DMs with certain off-the-shelf FMs.
For the first scenario, FATE-LLM (Fan et al., 2023) proposed a FedCoLLM (Federated Co-tuning LLM) framework (Fan et al., 2023) to co-optimize the server’s FM and clients’ DMs simultaneously through mutual distillation. Figure 11(a) illustrates the workflow of FedCoLLM. In FedCoLLM, clients’ DMs are initialized with an off-the-shelf FM (i.e., LLaMA-7B (Touvron et al., 2023)). In each communication round, clients fine-tune their DMs using local data and then send them to the server with secure aggregation (Bonawitz et al., 2016). The server, in turn, distills the knowledge mutually between its hosted FM (i.e., LLaMa-65B) and the global DM aggregated from clients’ local DMs to co-optimize both models. Subsequently, the server dispatches the global DM to all clients for further training.  Deng et al. (2023) proposed CrossLM that optimizes the server’s FM and clients’ DMs through synthetic data that are generated by the server’s FM and validated by clients’ DMs, as illustrated in Figure 11(b). Specifically, in each communication round, clients train their DMs using local data and then send them to the server asynchronously. The server, in turn, generates synthetic data and leverages received DMs to validate the quality of synthetic data. Subsequently, the server utilizes validation feedback from DMs and the validated synthetic data to fine-tune the server’s FM. Next, the server uses the validated synthetic data to fine-tune each client’s DM and dispatches the fine-tuned DMs to corresponding clients for further training. This concerted effort is aimed at co-optimizing both the FM and DMs. FedCoLLM and CrossLM adhere to the objective formulated in Eq.(4) with $K>1$, $\sum^{K}_{k=1}p^{C_{k}}=1$, and $p^{S}=0$. FedCoLLM has $\eta>0$ whereas CrossLM has $\eta=0$. The specific form of the loss $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ for co-optimizing the server’s $\mathcal{W}^{S}_{\text{FM}}$ and a client $k$’s $\mathcal{W}^{C_{k}}$ is formulated in Eq.(12).
@@
Figure 12. Federated parameter-efficient-finetuning of clients’ domain-specific FMs. Clients’ DMs are initialized with FMs and then learn from each other through federated fine-tuning of their PEFT modules (e.g., adapter, prompt model, and LoRA model). The server hosts no model and acts as an aggregation function.
For the second scenario, FedNLP (Lin et al., 2022) provided a benchmarking framework for evaluating federated learning algorithms (e.g., FedAvg, FedProx, and FedOPT) on common NLP tasks with FMs (e.g., BART). Woisetschläger et al. (2023b) provided a micro-level hardware benchmark for measuring Energy, Computational, and Communication efficiency when fine-tuning the FLAN-T5 model family, ranging from 80M to 3B parameters, using FL. Fine-tuning or training clients’ full FMs is computationally expensive, primarily because of the large sizes of FMs (Cai et al., 2023b). To alleviate this issue, various Parameter-Efficient Fine-Tuning (PEFT) methods have been proposed. As illustrated in Figure 12, these PEFT-based FL methods typically involve clients whose DMs are initialized with pre-trained FMs. During training, each client only uploads a PEFT module to the server for aggregation. The PEFT module can be an adapter, LoRA model, or prompt model that contributes to a small portion of the parameters of the original FM, thereby saving computational and communication overheads. These PEFT-based FL methods correspond to a special case of the objective formulated in Eq.(5) with $K>1$ and $\sum^{K}_{k=1}p^{C_{k}}=1$. The specific form of the loss $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ for fine-tuning client $k$’s $\mathcal{W}^{C_{k}}$ with PEFT methods is formulated in Eq.(13).
@@
As for now, few works following Eq.(5) explicitly consider protecting data privacy (i.e., $\eta>0$). FedPETuning (Zhang et al., 2023d), PEU+LoRA (Xu et al., 2023a), and FedPrompt (Zhao et al., 2023b) are representative ones. FedPETuning leverages four PEFT techniques, namely adapter tuning, prefix tuning, LoRA (Hu et al., 2022), and BitFit (Ben Zaken et al., 2022), to fine-tune clients’ DMs (i.e., $\text{Roberta}_{\text{base}}$) and demonstrated that federated learning combined with LoRA achieved the best privacy-preserving results. PEU+LoRA applies DP with Partial Embedding Updates (PEU) and LoRA to fine-tune clients’ DMs, achieving better privacy-utility-resource trade-offs than baselines. Zhao et al. (2023b) proposed FedPrompt that leverages federated learning combined with soft prompt tuning to fine-tune clients’ DMs (e.g., T5 (Raffel et al., 2020)), aiming to reduce communication cost while maintaining model accuracy. FedPrompt adopts local DP to protect the privacy of clients’ data, demonstrating there is a trade-off between privacy and model accuracy.
@@
Most methods following Eq.(5) focus on improving computation and communication efficiency (i.e., $\eta=0$). Chen et al. (2024) proposed FedDAT, in which each client fine-tunes its local multi-modal model (e.g., VAuLT (Chochlakis et al., 2022)) by leveraging a Dual-Adapter Teacher (DAT) module to achieve better convergence rate than using conventional PEFT methods. Li et al. (2022b) proposed FedSplitBERT, which splits the BERT encoder into a local part and a global part and shares only a quantized version of the global part with the server for aggregation, aiming to reduce the communication cost with minimal performance loss. FedIT (Zhang et al., 2023c) was proposed to leverage federated learning combined with LoRA to efficiently utilize clients’ diverse instructions stored on local devices to improve the performance of global FM (e.g., Vicuna (Chiang et al., 2023)). Che et al. (2023) proposed a parameter-efficient prompt tuning approach with adaptive optimization, named FedPepTAO, to fine-tune clients’ large language models (e.g., LLaMA-7B) in a non-IID setting. Halbe et al. (2023) proposed HePCo to address a continual federated learning problem. In HePCo, each client is deployed with a ViT model, and all clients share only a prompt model and a classifier (on top of the ViT) with the server to save communication costs. To tackle catastrophic forgetting brought by continual learning and data heterogeneity across clients, the server trains a generator based on current and previous tasks and distills the knowledge from the generator to the global prompt model and classifier at each communication round.  Wu et al. (2023b) proposed Fed-LPFM, in which each client uses its local data and a set of private FMs to train a proxy model, and clients’ proxy models are shared with and aggregated by the server. At inference, only the proxy models are used. As a result, the inference latency can be considerably reduced, and the data distribution gap between clients can be mitigated. Shi et al. (2024) proposed CLIP2FL that leverages an off-the-shelf CLIP model to assist federated learning on heterogeneous and long-tailed data. In each communication round, in addition to local training, each client distills the knowledge from a ViT-B/32 model to its local model, which consists of a feature extractor and a classifier and has a much smaller size than ViT-B/32, while the server generates federated features based on aggregated gradients to fine-tune a global classifier for all clients.
@@
Besides the aforementioned works, a potentially more demanding and intriguing avenue of research involves empowering edge devices (e.g., smartphones, wearables, and Internet of Things devices) with the knowledge and capabilities of FMs, thereby greatly augmenting the intelligence of systems facilitated by these devices. Take, for instance, the case of smart home systems (King et al., 2023) enhanced by FMs, which can elevate the quality of life for individuals, or the case of surveillance systems (Yao et al., 2019) fortified by FMs, leading to heightened efficiency and accuracy in detecting moving objects such as people and traffic. To achieve these visions, the primary challenge lies in effectively and efficiently adapting FMs to edge devices that have limited memory and computing resources. Federated learning presents a promising solution to this challenge, and several studies have investigated this direction. FedAdapter (Cai et al., 2023b) was proposed to improve the computational and communication efficiency of fine-tuning local FMs (e.g., BERT) resided in devices through PEFT (i.e., adapter). To achieve this, it dynamically adjusts the depth and width of adapters during training. FeS (Cai et al., 2023a) was proposed to address the problem of data scarcity in mobile NLP applications by utilizing pseudo-labeling. It employs a curriculum pacing mechanism to gradually speed up the pseudo-labeling speed, a representational filtering mechanism to pseudo-label samples that are most beneficial to learn from, and a co-planing mechanism to control the layer depth and capacity to be trained. Furthermore, FeS fine-tunes only bias parameters of FMs (e.g., RoBERTa-large) while keeping the other parameters frozen to improve training efficiency further. FwdLLM (Xu et al., 2023b) aims to enhance the efficiency of PEFT techniques. It observed that PEFT techniques mainly alleviate network bottlenecks but do not significantly improve training efficiency as they require backpropagation across the entire model to obtain gradients. To address this challenge, FwdLLM employs a backward propagation (BP)-free method to estimate gradients during the forward pass, thereby saving substantial memory and training time. To tackle system heterogeneity, Cho et al. (2023) proposed a heterogeneous LoRA (HLoRA) approach that can apply different rank LoRA modules to devices with heterogeneous and constrained computing resources by leveraging zero-padding and truncation for the aggregation. Woisetschläger et al. (2023a) conducted an in-depth study on hardware performance optimization for federated fine-tuning large language models at the edge. More specifically, It investigated the state-of-the-art deep learning accelerators as well as optimization levers for FL workloads on embedded hardware.